# An Approximation Algorithm for Monotone Submodular Cost Allocation

### v1

@@ Appendix A Convex program based on the Lovász extension
In this section, we show that LP-Rel is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5]. Let $N:=\{1,\ldots,n\}$. Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$. Recall that LP-Rel is defined as follows:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel can be translated into a feasible solution $(y_{1},\ldots,y_{k})$ to LP-Rel that has the same objective value and whose supports are chains. Take any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel. We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$. We first relabel the elements in $N$ so that $x^{i}_{1}\geq x^{i}_{2}\geq\cdots\geq x^{i}_{n}$. Let $x^{i}_{n+1}:=0$ and $S_{j}:=\{1,2,\ldots,j\}$ for each $j\in[n]$. Then, for all $S\subseteq N$ define $y_{i}(S)$ as follows:
@@
Then, $(y_{1},\ldots,y_{k})$ is a feasible solution to LP-Rel because for each $i\in[k]$ we have
@@
which implies that $(y_{1},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$. Similarly, we can translate any feasible solution to LP-Rel whose supports are chains into a feasible solution to LE-Rel that has the same objective value.
Since LE-Rel can be solved in polynomial time via the ellipsoid method [5], an optimal solution to LP-Rel whose supports are chains can be obtained in polynomial time. Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal. We note that LP-Rel is equivalent to LE-Rel-Primal.
@@ Appendix B $k$-polymatroid intersection
The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$. When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of LP-Rel. For the case when $f_{1},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].

### v2

@@ Appendix A Convex program based on the Lovász extension
In this section, we show that (LP) is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5]. Let $N:=\{1,\ldots,n\}$. Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$. Recall that (LP) is defined as follows:
$\textsc{(LP)}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel) can be translated into a feasible solution $(y_{1},y_{2},\ldots,y_{k})$ to (LP) that has the same objective value and whose supports are chains. Take any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel). We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$. We first relabel the elements in $N$ so that $x^{i}_{1}\geq x^{i}_{2}\geq\cdots\geq x^{i}_{n}$. Let $x^{i}_{n+1}:=0$ and $S_{j}:=\{1,2,\ldots,j\}$ for each $j\in[n]$. Then, for all $S\subseteq N$ define $y_{i}(S)$ as follows:
@@
Then, $(y_{1},y_{2},\ldots,y_{k})$ is a feasible solution to (LP) because for each $i\in[k]$ we have
@@
which implies that $(y_{1},y_{2},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$. Similarly, we can translate any feasible solution to (LP) whose supports are chains into a feasible solution to (LE-Rel) that has the same objective value.
Since (LE-Rel) can be solved in polynomial time via the ellipsoid method [5], an optimal solution to (LP) whose supports are chains can be obtained in polynomial time. Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal. We note that (LP) is equivalent to LE-Rel-Primal.
@@ Appendix B $k$-polymatroid intersection
The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$. When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of (LP). For the case when $f_{1},f_{2},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].